OpenStudy (anonymous):

Find the slope of the line passing through the points A(8, 7) and B(-2, -6)

OpenStudy (anonymous):

the slope is found by m=(y1-y2)/(x1-x2)=13/10

OpenStudy (saifoo.khan):

(-6-7)/(-2-8)

OpenStudy (saifoo.khan):

13/10

OpenStudy (anonymous):

the answer is 13/10